OpenStudy (anonymous):

in the sunshine, an upright pole 12 feet tall is casting a shadow 8 feet long. at the same time a nearby upright pole is casting a shadow 10 feet long. if the lengths of the shadows are proportional to the heights of the poles, what is the height in feet of the taller pole??

OpenStudy (anonymous):

solve the ratio \[\frac{8}{12}=\frac{x}{10}\]

OpenStudy (anonymous):

one step does it \[x=\frac{10\times 8}{12}\]

OpenStudy (anonymous):

ok

OpenStudy (nilankshi):

good

OpenStudy (anonymous):

but how do i know that isnt a triangle ques???

OpenStudy (anonymous):

is it because of the word proportional??

OpenStudy (amistre64):

similar triangles is the key

OpenStudy (anonymous):

so it is a triangle ques

OpenStudy (amistre64):

it is, but since they are similar triangles they have the same ration of parts

OpenStudy (anonymous):

so can u break down the equation for me please

OpenStudy (amistre64):

where h = height and s = shadow \[\frac{h_1}{s_1}=\frac{h_2}{s_2}\] if you know 3 of the elements; you can find the missing piece

OpenStudy (anonymous):

Be sure to place like elements in the same spots or your porportion will be wrong

OpenStudy (anonymous):

so 8x=120 then120/8 then 15

OpenStudy (amistre64):

its a good thing I named them similarly then lol ;)

OpenStudy (amistre64):

12 tall with a shadow 8 feet long x tall with a shadow 10 feet long 12/8 = x/10 120/8 = x \ looks good

OpenStudy (amistre64):

similar triangles is the key ...

OpenStudy (anonymous):

thanks for the help..!!